OpenStudy (anonymous):

Why is (3, –5) not a solution to the system of equations given below? {-x+4y=-15 {-2x-3y=-8 A. It satisfies the top equation but not the bottom equation. B. It satisfies the bottom equation but not the top equation. C. It doesn’t satisfy either the top or the bottom equations. D. It satisfies both the top and the bottom equations.

OpenStudy (ybarrap):

$$ -x+4y=-15 \\ -2x-3y=-8 $$ add -2 times the 1st equation to the 2nd: $$ -8y-3y=30-8=22\\ y=\cfrac{22}{-11}=-2\\ $$ From which you can get x=7 So (3, –5) doesn’t satisfy either the top or the bottom equations